# Explainability of Large Language Models using SMILE: Statistical Model-agnostic Interpretability with Local Explanations

### v1

@@ 4.3 Developing the Interpretable Surrogate Model
We are innovative in mapping the output space of the text generation model to a one-dimensional distance scale, accompanied by corresponding perturbation texts for the calculated distances [117]. This method enables us to train an interpretable model on the mapped output space of the text generation model, effectively demonstrating how input words influence text generation.
In our linear regression model, the vectors of the perturbation texts are critical, as they are treated as independent variables. Following the extraction of text embeddings, the distance between the outputs of these perturbation texts and the corresponding original text output is regarded as the response variable. Additionally [118], the degree of similarity between each perturbation text and the original text is assumed to represent the weight of the samples [102]. Eq. 16 outlines the square loss objective function for interpretable weighted linear regression:
$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=$ (16)
$\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\times\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$
Here, $\phi_{\text{gen}}$ represents the text generation model, and $\delta_{\text{WMD}}$ is the Word Mover’s Distance function applied to compute similarity between generated outputs. The weight $\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ is derived using a Gaussian kernel function and reflects the degree of similarity between the original prompt and each perturbed prompt. The term $\hat{f}_{t}(p_{\text{pert}}^{i})$ is a predicted value from the linear regression model.
After feature extraction, the $\hat{f}_{i}(p_{\text{pert}}^{i,01})$ function is a linear regression model trained over the one-hot vector space of perturbation texts $(p_{\text{pert}}^{i,01})$ and the calculated distances [102].
@@ 4.4.2 Stability
Stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [123]. In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations. Stability is a key property for fostering trust in explainable artificial intelligence (XAI) systems [14]. It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [15].
@@ 4.4.4 Fidelity
Fidelity is a metric to evaluate how well an explainable model aligns with a black-box model. The predictions of the black-box model, denoted as $f(x_{i})$, and the explainable model, $g(x_{i})$, for the $i$-th data perturbation are compared. Various loss functions and coefficient metrics quantify the similarity between the models’ predictions, measuring how effectively the explainable model approximates the black-box model.
The coefficient of determination $R^{2}$ is an accuracy-based metric that measures how well the explainable model correlates with the black-box model [124], and is defined in Eq. 18:
$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f}(X_{i}))^{2}}$ (18)
where $\bar{f}(X_{i})$ represents the average of $f(X_{i})$. Values of $R^{2}$ close to 1 indicate a strong alignment between the two models.
For scenarios where weighting is necessary, the weighted coefficient of determination $R_{w}^{2}$ provides an alternative version [16, 125], defined as Eq. 19:
$R_{w}^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f_{w}}(X_{i}))^{2}}$ (19)
where $\bar{f_{w}}(X_{i})$ represents the weighted average of $f(X_{i})$. These accuracy-based metrics, $R^{2}$ and $R_{w}^{2}$, offer insight into how well the explainable model approximates the black-box model.
To account for both sample size and the number of features used in the local explanation, the weighted adjusted coefficient of determination $\hat{R}_{w}^{2}$ [16] is used, as shown in Eq. 20:
@@
To measure the error or risk between the models’ predictions, error-based metrics, including the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE), are used [126]. These metrics quantify the weighted differences between the predictions of the explainable model and those of the black-box model. Lower values indicate better alignment. WMSE and WMAE are defined in Eqs. 21 and 22:
@@
Here, $w_{i}$ denotes the weight for each data point $i$, $f(X_{i})$ represents the predictions from the black-box model, $g(X_{i})$ represents the predictions from the explainable model, and $N_{p}$ is the total number of data points.
Additionally, the mean $L_{1}$ and mean $L_{2}$ losses, defined below, are commonly used to assess prediction discrepancies [127, 16], as shown in Eqs. 23 and 24:
@@
These error-based metrics, including MSE, MAE, $L_{1}$, and $L_{2}$, quantify the differences between the black-box and explainable models’ predicted scores, reflecting the risk of prediction discrepancies.
We compute fidelity across various scenarios. Fidelity is measured by comparing the predictions of the explainable model to those of the black-box text generation model. We analyse how different perturbations in the input text affect fidelity scores and how various distance metrics for text-to-text comparisons perform when using linear regression and Bayesian ridge as surrogate models.
@@ 5 Experimental Results
Experiments were conducted using a range of datasets and four state-of-the-art text generation models, OpenAI GPT [1], LLaMA [5], and Claude-AI [128] to ensure a comprehensive assessment.
@@ 5.1.2 Reasoning Path and Token Attribution
Beyond token-level importance visualisations, the proposed framework provides structured reasoning paths that illustrate how the model logically connects input elements to produce its response. Figure 8 presents an example where the SMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram. This combined visualisation offers a holistic understanding of local word-level contributions and the logical structure of the model’s decision-making.
Such comprehensive visual explanations align with advanced techniques used by leading AI research organisations, including Anthropic and OpenAI, while providing greater flexibility across diverse tasks and models. As shown in Figure 8, SMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [129].
Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the SMILE framework. The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates SMILE’s token-level importance heatmap for the same prompt. [129]
5.1.3 Word-Level Attribution Analysis on MMLU Using SMILE
To evaluate the interpretability of ChatGPT using the SMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [130]. This benchmark contains four-choice multiple-choice questions covering various academic and professional subjects, including science, history, law, and the humanities. For this study, we accessed a pre-formatted version of the dataset available on Kaggle [131] to facilitate prompt construction and experimentation.
Figure 9 shows a set of attribution heatmaps that visualise which parts of the input prompt most influenced the model’s decision. Darker red tones indicate higher influence, while neutral or irrelevant tokens are shaded lighter. These visualisations demonstrate SMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
Figure 9: Attribution heatmaps generated by SMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
@@ 5.2 Quantitative Results
In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to asses our explainability method across different text generation models and different scenarios.
@@ 5.2.2 Stability
To quantify stability, we use the Jaccard index, a metric that measures the similarity between sets by comparing their overlap. For text generation models such as OpenAI GPT, LLAMA, and Claude-AI, we tested ten different prompts and extracted the coefficients and weights assigned to each word. To evaluate stability, we added the token ‘‘ at the end of each prompt as input text, extracted the coefficients again, and computed the Jaccard index to compare the sets of coefficients before and after the perturbation. Finally, we calculated the average Jaccard index for each model, and the results are presented in Table 2.
Table 2: Stability across different models for ten prompts and 30 perturbations
@@ 5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
[…]
@@ 5.2.7 Computation complexity
We apply our model to three instruction-based text generation frameworks: OpenAI GPT, LLaMA, and Claude-AI. We also evaluate the execution time of different explainability methods: LIME, SMILE, and Bay-LIME. Each method is tested with 60 perturbations on the same hardware to ensure consistency in computational conditions.
The models exhibit notable differences in execution time, primarily due to how each explainability method computes similarity and assigns weights. SMILE, in particular, employs the Wasserstein distance to quantify text dissimilarities, making it substantially more computationally expensive than both LIME and Bay-LIME. As reported in Table 9, SMILE consistently incurs higher execution times across all three model frameworks. This pattern also holds across other frameworks, indicating that SMILE’s computational overhead is a systemic characteristic of its underlying methodology.
Wasserstein distance is known for capturing subtle differences between text embeddings, making it especially effective in high-dimensional semantic spaces [101]. However, this precision comes at a cost. Computing the Wasserstein-2 distance involves operations on covariance matrices with time complexity proportional to $O(d^{3})$, where $d$ is the embedding dimensionality. When applied across $N$ samples, the total complexity becomes $O(Nd^{3})$. In contrast, LIME relies on cosine similarity, which has a much lower computational cost of $O(Nd)$.
This disparity arises because Wasserstein-2 distance requires solving matrix optimisation problems over $d$-dimensional Gaussian distributions [132]. While this added complexity leads to more stable and expressive explanations, as demonstrated by SMILE, it significantly increases runtime. Conversely, LIME avoids this overhead by utilising simpler vector-based similarity measures.
In summary, SMILE offers greater fidelity and stability but at a higher computational cost. Its use of the Wasserstein distance makes it ideal for tasks where precision and interpretability are essential, even if that entails longer execution times. LIME and Bay-LIME, while faster, may trade off some robustness in favour of efficiency.
@@
Framework LIME SMILE Bay-LIME
@@
5.2.8 Optimised Instructions from Google Research
To further explore how instruction phrasing influences interpretability, we applied SMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [133]. These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
We first applied SMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem. ” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, SMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
Figure 10: Word-level importance in a high-performing instruction from Google Research, identified using SMILE with GPT-3.5.
Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using SMILE with GPT-4.

### v2

@@ 4.3 Developing the Interpretable Surrogate Model
To interpret the influence of individual words on text generation, we map the output space of the generation model to a one-dimensional distance scale, where each point corresponds to a perturbed prompt [128]. This enables the training of a local surrogate model that relates changes in input structure to variations in output behavior.
In this weighted linear regression framework, the perturbation vectors (i.e., binary or one-hot encodings of included or excluded tokens) serve as independent variables. The response variable is the Wasserstein distance between the output of a perturbed prompt and that of the original prompt. Each perturbation is assigned a weight based on its semantic similarity to the original prompt [114, 129].
The loss function minimized during training is a weighted squared error:
$\text{Loss}(\phi_{\text{gen}},\delta_{\text{WMD}},p_{\text{org}})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})\cdot\left(W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})-\hat{f}_{t}(p_{\text{pert}}^{i})\right)^{2}$ (16)
Here:
  • $\phi_{\text{gen}}$ is the text generation model,
  • $\delta_{\text{WMD}}$ denotes the function used to embed generated texts,
  • $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance between the generated outputs,
  • $\pi_{i}$ is the similarity-based kernel weight for the $i$-th perturbation,
  • and $\hat{f}_{t}(p_{\text{pert}}^{i})$ is the prediction from the linear surrogate model.
The surrogate model $\hat{f}_{t}$ is trained on the one-hot vector representation of the perturbation prompts $p_{\text{pert}}^{i,01}$, using the computed Wasserstein distances as targets [114].
@@ 4.4.2 Stability
Stability in an explainable model refers to its ability to provide consistent explanations even when minor changes are made to the input data [134]. In other words, small perturbations in the input should not significantly affect the model’s predictions or explanations. Stability is a key property for fostering trust in explainable artificial intelligence or XAI systems [61]. It reassures users that the model’s interpretations are not arbitrary or overly sensitive to irrelevant variations [10].
@@ 4.4.4 Fidelity
Fidelity is a metric used to evaluate how closely an explainable model replicates the behavior of a black-box model. Specifically, the predictions of the black-box model denoted as $f(X_{i})$, are compared to those of the explainable model, $g(X_{i})$, for each input perturbation $i$. A variety of accuracy-based and error-based metrics are employed to assess this alignment, thereby quantifying how well the explainable model approximates the responses of the black-box model.
Let $N_{p}$ denote the total number of perturbation samples (i.e., data points). One of the most commonly used accuracy metrics is the coefficient of determination $R^{2}$, which measures how well the explainable model’s predictions correlate with those of the black-box model [135]. It is defined as:
$R^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f})^{2}}$ (18)
where $\bar{f}$ represents the mean of the black-box model’s predictions $f(X_{i})$. A value of $R^{2}$ close to 1 indicates a strong alignment between the models.
In cases where data samples have different levels of importance, a weighted variant is used. The weighted coefficient of determination $R_{w}^{2}$ is given by [136, 137]:
$R_{w}^{2}=1-\frac{\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}}{\sum_{i=1}^{N_{p}}(f(X_{i})-\bar{f}_{w})^{2}}$ (19)
Here, $\bar{f}_{w}$ is the weighted mean of the black-box predictions. These accuracy-based metrics ($R^{2}$ and $R_{w}^{2}$) provide insight into how closely the explainable model aligns with the black-box model.
To account for both sample size and model complexity (i.e., number of features used), the weighted adjusted coefficient of determination $\hat{R}_{w}^{2}$ is used [136]:
@@
Where $N_{s}$ is the number of features in the interpretable model.
In addition to correlation-based metrics, error-based metrics are widely used to quantify differences in prediction values. The classical unweighted forms Mean Squared Error (MSE) and Mean Absolute Error (MAE) are:
$\text{MSE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(f(X_{i})-g(X_{i}))^{2}$ (21)
$\text{MAE}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|f(X_{i})-g(X_{i})|$ (22)
For situations requiring sample-level importance, the Weighted Mean Squared Error (WMSE) and Weighted Mean Absolute Error (WMAE) are defined as [138]:
@@
Here, $w_{i}$ denotes the sample weight associated with the $i$-th perturbation.
Moreover, the mean $L_{1}$ and mean $L_{2}$ norms are also applied to evaluate absolute and squared differences respectively [139, 136]:
@@
These metrics, MSE, MAE, WMSE, WMAE, $L_{1}$, and $L_{2}$—together quantify how accurately the explainable model reproduces the black-box model’s predictions and reflect the risk of prediction discrepancies across the perturbation space.
We compute fidelity across various scenarios by comparing the explainable model’s predictions to those of the black-box text generation model. This includes analyzing how different input perturbations affect fidelity scores and comparing the performance of various distance metrics and surrogate models, such as linear regression and Bayesian ridge regression, in aligning with the black-box behavior.
@@ 5 Experimental Results
Experiments were conducted using a range of datasets and three state-of-the-art text generation models, OpenAI GPT [1], LLaMA [19], and Claude-AI [140] to ensure a comprehensive assessment.
@@ 5.1.2 Reasoning Path and Token Attribution
Beyond token-level importance visualisations, the proposed framework provides structured reasoning paths that illustrate how the model logically connects input elements to produce its response. Figure 8 presents an example where the gSMILE method not only assigns attribution scores to individual tokens but also generates a reasoning flow diagram. This combined visualisation offers a holistic understanding of local word-level contributions and the logical structure of the model’s decision-making.
Such comprehensive visual explanations align with advanced techniques used by leading AI research organisations, including Anthropic and OpenAI, while providing greater flexibility across diverse tasks and models. As shown in Figure 8, gSMILE’s explainability aligns with or exceeds the interpretability achieved by leading industry methods [141].
Figure 8: Comparison between attribution visualisations by Anthropic’s Attribution Graphs and the gSMILE framework. The upper section shows the reasoning path identified by Anthropic’s method, while the lower section illustrates gSMILE’s token-level importance heatmap for the same prompt. [141]
5.1.3 Word-Level Attribution Analysis on MMLU Using gSMILE
To evaluate the interpretability of ChatGPT using the gSMILE framework, we used examples from the MMLU (Massive Multitask Language Understanding) dataset [142]. This benchmark contains four-choice multiple-choice questions covering various academic and professional subjects, including science, history, law, and the humanities. For this study, we accessed a pre-formatted version of the dataset available on Kaggle [143] to facilitate prompt construction and experimentation.
Figure 9 shows a set of attribution heatmaps that visualise which parts of the input prompt most influenced the model’s decision. Darker red tones indicate higher influence, while neutral or irrelevant tokens are shaded lighter. These visualisations demonstrate gSMILE’s ability to reveal the model’s internal reasoning by tracing the most impactful words in multi-choice question contexts.
Figure 9: Attribution heatmaps generated by gSMILE using ChatGPT (GPT-4) on selected samples from the MMLU dataset. Highlighted words represent those with the highest influence on the model’s output.
5.1.4 Optimised Instructions from Google Research
To further explore how instruction phrasing influences interpretability, we applied gSMILE to analyse two high-performing prompts introduced by Google Research through their OPRO framework [144]. These prompts were designed to enhance reasoning in LLMs and were explicitly optimised for zero-shot performance on mathematical tasks.
We first applied gSMILE to GPT-3.5 using 1024 perturbations on the instruction: “A small quantity of arithmetic and a logical approach will help us quickly solve this problem. ” As shown in Figure 10, the most influential words identified were “arithmetic” and “approach,” and “solution”, confirming their importance in shaping reasoning-oriented outputs.
We then extended the analysis to GPT-4 using the instruction: “Let’s combine our numerical command and clear thinking to quickly and accurately decipher the answer.” As illustrated in Figure 11, gSMILE revealed that “numerical,” “command,” and “answer” had the highest positive influence.
Figure 10: Word-level importance in a high-performing instruction from Google Research, identified using gSMILE with GPT-3.5.
Figure 11: Word-level importance in an instruction optimized by Google Research, analysed using gSMILE with GPT-4.
@@ 5.2 Quantitative Results
In this part, we apply some evaluation metrics, including accuracy, stability, consistency, fidelity, and computation complexity, to assess our explainability method across different text generation models and different scenarios.
@@ 5.2.1 Accuracy
Among the evaluated models, Claude-AI demonstrated the strongest performance, achieving the highest scores in all three metrics: 0.82 in ATT ACC, 0.67 in ATT F1, and 0.88 in ATT AUROC. These results indicate Claude-AI’s superior ability to consistently focus on essential input tokens across varying prompts. In contrast, LLAMA achieved the second-highest accuracy but showed a notably lower F1-score (0.40), suggesting potential inconsistency in identifying key tokens. OpenAI GPT yielded balanced results but lagged slightly behind Claude-AI in all metrics, highlighting room for improvement in its attention fidelity. Ll, Claude-AI provides more reliable and semantically focused attention compared to its counterparts.
@@ 5.2.2 Stability
To quantify stability, we use the Jaccard index, a metric that measures the similarity between sets by comparing their overlap. For text generation models such as OpenAI GPT, LLAMA, and Claude-AI, we tested ten different prompts and extracted the coefficients and weights assigned to each word. To evaluate stability, we added only the token  at the end of the text, extracted the coefficients again, and computed the Jaccard index between the coefficients before and after the perturbation. The averaged Jaccard indices across all ten prompts for each model are reported in Table 2. Higher scores indicate greater stability under minor input changes.
Table 2: Stability scores across models based on average Jaccard index over ten prompts
@@ 5.2.3 Consistency
Among all models, OpenAI-GPT exhibited the highest consistency, achieving a variance of 0.0000 and a standard deviation of 0.0046. This indicates that its predictions are the most stable across repeated runs, confirming its robustness and reliability in generating interpretable outputs.
@@ 5.2.4 Fidelity for Different Instruct image editing diffusion based Models (IED)
From the results, we observe that Claude-AI achieves the lowest WMSE, WMAE, and mean-L1, indicating superior performance in minimizing error. On the other hand, OpenAI-GPT performs best in terms of $R^{2}_{\omega}$, mean-L2, and adjusted $R^{2}_{\hat{\omega}}$, demonstrating stronger alignment and generalization of the surrogate model. LLaMA ranks in the middle across most metrics but does not lead in any. This suggests that different models excel in different fidelity dimensions, highlighting the importance of multi-metric evaluation when interpreting black-box model behavior.
@@ 5.2.5 Fidelity Across Different number of Text Perturbations
For OpenAI GPT and LLaMA, increasing the number of perturbations generally improves fidelity across most metrics, with 256 perturbations achieving the best WMSE, WMAE, and loss values. Interestingly, for Claude-AI, 64 perturbations yield the lowest WMSE and WMAE, while 256 achieves the best loss (mean-L2 and mean-L1). However, the highest $R^{2}$ scores for all three models occur at 32 perturbations, suggesting that beyond a certain point, more perturbations may dilute interpretability despite improving local fit. Therefore, there is a trade-off between the number of perturbations and the reliability of the surrogate model, depending on the metric of interest.
@@ 5.2.6 Fidelity for Different Distance Metrics and Surrogate Models
Fidelity is also computed by comparing various distance metrics for text generation models such as OpenAI GPT using text-to-text comparisons. The surrogate models used are Weighted Linear Regression (WLR) and Bayesian Ridge (BayLIME), as shown in Table 8. We explore different combinations of distance metrics, including Cosine similarity and Wasserstein Distance (WD). Fidelity is measured using WMSE, WMAE, mean-$L_{1}$, mean-$L_{2}$, and coefficient of determination ($R^{2}_{\omega}$, $R^{2}_{\hat{\omega}}$).
@@
Our Proposed Method: gSMILE.
Based on the results in Table 8, the combination where Wasserstein Distance (WD) is used for both input similarity and output shift consistently yields the highest values of the weighted coefficient of determination ($R^{2}_{\omega}$ and $R^{2}_{\hat{\omega}}$). Specifically, underweighted linear regression, the WD–WD configuration achieves $R^{2}_{\omega}=\textbf{0.7104}$ and $R^{2}_{\hat{\omega}}=\textbf{0.6409}$, which significantly outperform other configurations.
Motivated by these findings, we introduce our method, called gSMILE (generalized SMILE), which utilizes the Wasserstein distance for both input and output comparisons. This choice enhances the surrogate model’s fidelity and stability in capturing local shifts in model behavior.
We use gSMILE throughout subsequent sections as the foundation for explainability evaluations on large language models (LLMs). These include assessments of fidelity, consistency, and stability, ensuring the proposed approach is both theoretically grounded and empirically adequate.
@@ 5.2.7 Computation Complexity
We evaluate the computational efficiency of our proposed method, gSMILE, using three instruction-based text generation models: OpenAI GPT, LLaMA, and Claude-AI. Although these models differ in scale and architecture, the objective here is not to compare the models themselves but rather to assess the performance of different explainability methods applied to each model under identical hardware and runtime conditions. All evaluations are conducted with 60 perturbations per input prompt.
We compare three explainability approaches—LIME, gSMILE, and Bay-LIME—based on their execution time. As reported in Table 9, gSMILE consistently requires the highest runtime across all model frameworks. This is mainly due to its reliance on the Wasserstein distance to compute similarity between output distributions, which involves solving high-dimensional optimization problems with time complexity $O(Nd^{3})$, where $N$ is the number of perturbations and $d$ is the embedding dimensionality [145]. By contrast, LIME and Bay-LIME utilize cosine similarity with lower complexity $O(Nd)$, making them faster but potentially less precise.
Despite the higher computational cost, gSMILE delivers significantly better fidelity, stability, and robustness to adversarial perturbations—essential properties for trustworthy explanations. These advantages are supported by results in Tables 8 and 2 and align with findings from prior work on SMILE [11], which demonstrated superior robustness against adversarial input variations.
These findings suggest that gSMILE’s computational overhead is justified by the quality of explanations it provides, making it highly suitable for applications where interpretability, consistency, and resistance to manipulation are paramount.
@@
Framework LIME gSMILE Bay-LIME